40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 01:58:57 2020, gautam, Update, Electronics, Some electronics thoughts IMCsensingNoise.pdfIMCsensingNoise.pdf
    Reply  Wed Jun 3 01:40:13 2020, gautam, Update, Electronics, More electronics woes glitchySeis2.pngglitchySeis3.png
Message ID: 15365     Entry time: Wed Jun 3 01:40:13 2020     In reply to: 15347
Author: gautam 
Type: Update 
Category: Electronics 
Subject: More electronics woes 

There were many locklosses from the point where the arm powers were somewhat stabilized. Attachments #1 and #2 show two individual locklosses. I think what is happening here is that the BS seismometer X channel is glitching, and creating a transient in the angular feedforward filter that blows the lock. The POP QPD based feedback loop cannot suppress this transient, apparently. For now, I get around this problem by boosting the POP QPD feedback loop a bit, and then turning the feedforward filters off. The fact that the other seismometer channels don't report any transient makes me think the problem is either with the seismometer itself, or the readout electronics. The seismometer masses were recently recentered, so I'm leaning towards the latter.

I didn't explicitly check the data, but I am reasonably certain the same effect is responsible for many PRMI locklosses even with the arms held off resonance (though the tolerance to excursions there is higher). Pity really, the feedforward filters were a big help in the lock acquisition...

Attachment 1: glitchySeis2.png  260 kB  Uploaded Wed Jun 3 02:45:31 2020  | Hide | Hide all
glitchySeis2.png
Attachment 2: glitchySeis3.png  282 kB  Uploaded Wed Jun 3 02:45:38 2020  | Hide | Hide all
glitchySeis3.png
ELOG V3.1.3-